# Cascaded Self-Evaluation Augmented Training for Efficient Multimodal Large Language Models

### v1

@@
Abstract
Efficient Multimodal Large Language Models (EMLLMs) have witnessed rapid development in recent years. Introducing Chain-of-Thought (CoT) during inference and leveraging self-evaluation step-by-step have further enhanced their performance. However, due to limited amount of the parameters, EMLLMs often perform poorly when directly using self-evaluation during inference. How to synthesize evaluation data? How much data to synthesize? What are the optimal training and inference strategies? What prompts to use? These are unexplored yet essential questions that need to be addressed to enhance the self-evaluation capabilities of EMLLMs.
To this end, we propose a method called Self-Evaluation Augmented Training (SEAT). We utilize more powerful EMLLMs to perform CoT reasoning, select beneficial data, and generate evaluations. The synthesized data is then used to train the EMLLMs. However, EMLLMs can’t handle long prompt inputs, and long CoT inference and self-evaluation outputs. Furthermore, synthesized data significantly diminishes its original CoT reasoning capabilities. To overcome these limitations, we further introduce Cascaded Self-Evaluation Augmented Training (Cas-SEAT). Cas-SEAT splits the long integrated prompts, which combine CoT reasoning and self-evaluation functionalities, into shorter cascaded prompts, each dedicated to one of these tasks. Additionally, we minimize the cost of Cas-SEAT to ensure its applicability in resource-constrained environments. During the synthesis of evaluation data, we use open-source EMLLMs with only 7B parameters and annotate a small amount of data using short prompts. Experiments show that, compared to existing methods, Cas-SEAT significantly enhances the self-evaluation capabilities of EMLLMs. It improves 19.68%, 55.57%, and 46.79% on the MathVista, Math-V, and We-Math datasets, respectively. Additionally, as the first work to explore self-evaluation augmented training for EMLLMs, the Cas-SEAT Dataset we constructe to enhance the self-evaluation capabilities of EMLLMs also provides valuable resources and references for future research in this field.
1 Introduction
Figure 1: Several modes of self-evaluation: (a) A multimodal dataset containing images, questions, and answers. (b) The typical approach in existing methods, where the model performs self-evaluation directly during inference. (c) The primary model synthesizes self-evaluation data and uses it for its own training. (d) A more powerful model evaluates the primary model’s outputs, and the primary model learns from this feedback to develop self-evaluation capabilities. (e) A more powerful model evaluates a small subset of the primary model’s outputs, combines the evaluations with CoT data, and feeds them into the primary model. This approach also decomposes prompts and tasks to enable cascading self-evaluation capabilities.
@@ 1 Introduction
In recent years, multimodal large language models (MLLMs) Liu et al. (2024); Wang et al. (2024) have advanced rapidly, achieving remarkable performance. However, broader training and deployment demands often come from resource-constrained settings, such as universities, hospitals, and communities. In these scenarios, the application of larger models, such as 72B or 40B-parameter MLLMs, faces significant limitations. This has spurred the emergence of efficient multimodal large language models (EMLLMs) with smaller parameter sizes, such as 7B, 2B, or even 1B. Today, EMLLMs are approaching the performance of early 40B-parameter MLLMs, thanks to advancements in model architecture, training methods, and techniques like Chain-of-Thought (CoT) Wei et al. (2022c). CoT, in particular, enables EMLLMs to explicitly model intermediate reasoning steps, producing more interpretable, logically coherent, and contextually relevant responses across a wide range of tasks involving text, images, and other modalities.
Despite these advancements, the further improvement of EMLLMs faces two major bottlenecks. First, the CoT data available for training is insufficient in both quantity and quality. Existing data, often sourced from web crawls or generated by models without strict quality control, suffers from inconsistencies, incomplete reasoning steps, logical gaps, and even incorrect reasoning. Such flawed samples not only fail to enhance model performance but may also degrade it. Second, while recent research suggests that self-evaluation can improve the accuracy and reliability of reasoning, these efforts primarily focus on refining prompts during the inference stage Xu et al. (2024); Lightman et al. (2024). They guide models to evaluate each reasoning step for accuracy but rarely generate evaluation data during training to enhance the self-evaluation capabilities of LLMs and MLLMs. Research on EMLLMs in this area remains non-existent. Due to their smaller parameter sizes, EMLLMs struggle to evaluate their own reasoning steps effectively or correct errors (as evidenced by the negligible performance improvement after evaluation in Figure 3’s black bar).
To address these challenges, we propose Self-Evaluation Augmented Training (SEAT). SEAT leverages more powerful EMLLMs to perform CoT reasoning and evaluate it. By refining these evaluations, we generate high-quality reasoning samples that help improve the overall reasoning quality and self-evaluation capabilities of the target model (referred to as the "main model" in this paper). However, the limited parameter scale of EMLLMs introduces two challenges to this process: (1) integrating reasoning and evaluation tasks into a single prompt leads to excessively long prompts during both input and output stages, overwhelming EMLLMs. For EMLLMs to understand that they need to perform CoT inference and self-evaluation, a long prompt is required. The output, consisting of step-by-step reasoning and the corresponding self-evaluation for each step, has to become even longer. (2) introducing extended evaluation data into EMLLMs can diminish their inherent CoT reasoning capabilities (as evidenced by SEAT’s poor inference and evaluation performance in Table  1, 2, 3, and 4).
To overcome these limitations, we further propose the Cascaded Self-Evaluation Augmented Training (Cas-SEAT) method. Cas-SEAT decouples the reasoning and self-evaluation processes into two separate tasks using two cascaded short prompts instead of a single lengthy integrated prompt. Regarding data handling, Cas-SEAT annotates only a small subset of data using the evaluation prompt in the evaluation phase and mixes this data with the reasoning prompt annotations from the reasoning phase. This approach ensures that EMLLMs first enhance their CoT reasoning capabilities before learning self-evaluation, as we believe that robust CoT reasoning is the foundation for effective self-evaluation. We also minimize the cost of Cas-SEAT to ensure its applicability in resource-constrained environments. During evaluation data synthesis, we use only open-source EMLLMs with 7B parameters and annotate a small dataset with short prompts. This method primarily generates high-quality reasoning samples, adding minimal self-evaluation data to optimize both data utility and training efficiency. Experimental results show that Cas-SEAT significantly improves self-evaluation capabilities compared to existing methods, with performance gains of 19.68%, 55.57%, and 46.79% on the MathVista, Math-V, and We-Math datasets, respectively. Equally important, as the first work to explore self-evaluation augmented training for EMLLMs, both Cas-SEAT and the Cas-SEAT Dataset we constructed provide a valuable foundation and reference for future research in this domain.
@@
  • We propose SEAT, exploring how EMLLMs improve their CoT reasoning and self-evaluation capabilities by learning from CoT-generated self-evaluation data. To the best of our knowledge, this is the first study of its kind on EMLLMs and is rare even among LLMs and MLLMs.
@@
  • We designed Cas-SEAT to effectively address the challenges EMLLMs face with handling long prompt inputs and lengthy CoT and self-evaluation outputs, successfully enhancing self-evaluation capabilities while maintaining CoT reasoning abilities.
@@
  • We construct the Cas-SEAT Dataset, the first dataset tailored for self-evaluation augmented training of EMLLMs. It is low-cost yet effective, facilitating future research.
  • We conducted extensive experiments across multiple datasets, exploring different model architectures (LLaVA-1.5, Qwen2-VL) and parameter scales (7B, 2B). We performed a detailed analysis of overall performance, variations in performance across different aspects, and further investigated the CoT reasoning ability and self-evaluation capability. The experimental results demonstrate that utilizing Cas-SEAT significantly enhances performance.
2 Related Work
CoT Reasoning.
@@ 2 Related Work
CoT reasoning Huang et al. (2022); Hoffmann et al. (2022); Chowdhery et al. (2022); Wei et al. (2022a); Zhang et al. (2022); Kojima et al. (2022) enhances the performance of LLMs by forcing them to perform step-by-step reasoning. The initial idea was proposed by Wei et al. (2022b), demonstrating that simply modifying the prompt could significantly improve model performance on complex tasks. Following Wei et al. (2022b), numerous studies have shown that LLMs can perform CoT reasoning through zero-shot CoT prompting (Kojima et al. (2022)) and few-shot CoT prompting (Wei et al. (2022b); Zhang et al. (2022)). Improvements have been made in areas such as self-consistency (Wang et al., 2022b), least-to-most prompting (Zhou et al., 2022), dynamic least-to-most prompting (Drozdov et al., 2022), bootstrapping training (Zelikman et al., 2022), and self-training (Huang et al., 2022). Wang et al. (2022b, a) refined the original CoT prompting by marginalizing over diverse reasoning paths, while Zelikman et al. (2022); Huang et al. (2022) improved CoT prompting through bootstrapping on self-generated CoT prompts. Li et al. (2022a) introduced a voting classifier to filter sampled CoT prompts before the final prediction. Additionally, prompt enhancement and selection (Shum et al., 2023), meta-heuristic approaches (Pan et al., 2023), and meta-graph prompting (Pan et al., 2024) have further advanced standard CoT prompting. Existing CoT studies pay little attention to the correctness of synthesized CoT data, especially lacking focus on the CoT capabilities of EMLLMs. CoT reasoning is more challenging for EMLLMs. Our Cas-SEAT improves the quality of synthesized training data through self-evaluation and enhances the CoT reasoning capabilities of EMLLMs.
LLM Self-Evaluation.
Recent research on LLM calibration has shown that current LLMs can achieve well-calibrated predictions in specific tasks (Rae et al., 2021; Kadavath et al., 2022; Guo et al., 2017; Kadavath et al., 2022; Jiang et al., 2021; Kuhn et al., 2023; Fu et al., 2023). For instance, Liu et al. Liu et al. (2023a) leveraged CoT prompting to use GPT-4 for text evaluation. Kocmi and Federmann Kocmi and Federmann (2023); Xu et al. Xu et al. (2023) designed methods focused on enabling LLMs to produce more fine-grained correction annotations. Additionally, Koo et al. Koo et al. (2023), Zheng et al. Zheng et al. (2023), Chang et al. Chang et al. (2023), Deutsch et al. Deutsch et al. (2022), and Liu et al. Liu et al. (2023b) explored issues of self-reinforcing bias and fairness in LLMs. It is worth noting that scaling up model size plays a crucial role in improving calibration (Rae et al., 2021; Wei et al., 2022a). As LLMs exhibit strong calibration capabilities, more research has focused on prompting LLMs for self-evaluation (Golovneva et al., 2022; Zhang et al., 2023; Shinn et al., 2023; Madaan et al., 2024; Paul et al., 2023). Although dedicated evaluators have proven effective in assessing the correctness of reasoning data (Li et al., 2022b; Xu et al., 2024), self-evaluation is more efficient compared to using external models for evaluation, as it eliminates the need for additional annotations and evaluators (Li et al., 2022a). Unlike existing methods that optimize the reasoning process through self-evaluation, our approach uses synthetic self-evaluation data to enhance model training. Moreover, we focus on EMLLMs, which introduce additional challenges to the training and reasoning processes of self-evaluation.
3 Methodology
@@
Figure 2: Overview of the method. It illustrates the prompts designed for Augmented Self-Evaluation and Augmented Cascading Self-Evaluation, along with the corresponding training data generated.
3.1 Problem Formulation and Notations
Data and Model.
@@ 3.1 Problem Formulation and Notations
The primary EMLLM model is denoted as $\mathcal{M}$, with parameters $\Theta_{0}$. To achieve better performance on the target task, $\mathcal{M}$ needs to be fine-tuned on the dataset $\mathcal{D}$. We use $\{\mathcal{X},\mathcal{Y}\}$ to represent a sample, $\mathcal{X}$ include images $\mathcal{I}$ and queries $\mathcal{Q}$, and $\mathcal{Y}$ represents answers. After fine-tuning, $\Theta_{0}$ will be optimized to $\Theta$. The output of $\mathcal{M}$ are denoted as $\hat{\mathcal{Y}}$. $\mathcal{D}_{\rm{se}}$ represents the self-evaluation data generated by $\mathcal{M}$. A more powerful EMLLM model is denoted as $\mathcal{G}$, which generates CoT reasoning data based on $\mathcal{D}$, denoted as $\mathcal{S}_{\rm{cot}}$. After evaluating $\mathcal{S}_{\rm{cot}}$, $\mathcal{G}$ produces self-evaluation training data, denoted as $\mathcal{S}_{\rm{se}}$. A small subset of $\mathcal{S}_{\rm{cot}}$ is selected, denoted as $\mathcal{S}_{\rm{cot}}^{\rm{sub}}$, and after being evaluated by $\mathcal{G}$, generates cascaded self-evaluation training data, denoted as $\mathcal{S}_{\rm{cse}}$.
Prompt.
@@
The prompts used for $\mathcal{M}$ inference and self-evaluation are denoted as $P$ and $P_{se}$, respectively. The prompts used by $\mathcal{G}$ to generate cascading CoT data and evaluation data are denoted as $P_{\rm{cot}}$ and $P_{se}$, respectively. The prompts used by $\mathcal{G}$ to generate CoT data and evaluation data are denoted as $P_{\rm{cot}}$ and $P_{cse}$, respectively. Identical symbols indicate identical prompts.
Formula.
We formalize the existing studies to facilitate a formula-based comparison of their differences. $l$ represents the loss function.
@@
Augmented CoT:
$&\mathcal{S}_{\rm{cot}}=\mathcal{G}(P_{\rm{cot}},\mathcal{X};\Theta)\\ &\operatorname*{arg\,min}_{\Theta}\mathcal{L}=\sum_{\mathcal{X},\mathcal{Y}\in\mathcal{S}_{\rm{cot}}}l(\mathcal{Y},\mathcal{M}(P\!\!\quad\text{or}\!\!\quad P_{\rm{cot}},\mathcal{X};\Theta))\end{aligned}\right.$
@@
3.2 SEAT
Figure 3: Comparison of evaluation in terms of model performance improvement

### v2

# Cascaded Self-Evaluation Augmented Training for Lightweight Multimodal LLMs

@@
Abstract.
Efficient Multimodal Large Language Models (EMLLMs) can improve performance through Chain-of-Thought (CoT) reasoning, but they have poor self-evaluation capabilities during the CoT reasoning process. This is due to their tendency to simplify the reasoning process and the degradation of self-evaluation ability during downstream task fine-tuning. To address this, we intuitively propose Self-Evaluation Augmented Training (SEAT), which uses more powerful EMLLMs to evaluate CoT reasoning data. The evaluation data is then used to train EMLLMs. However, due to the difficulties EMLLMs face with processing long token input-output sequences, and the degradation of self-evaluation ability as a basis for CoT reasoning, the SEAT method is not fully adapted. Therefore, we further propose Cascaded Self-Evaluation Augmented Training (Cas-SEAT), which converts long prompts into cascaded short prompts, each focusing on a specific task. Additionally, we mix CoT reasoning and self-evaluation data to preserve its CoT reasoning ability while enhancing the self-evaluation capability of EMLLMs. We also conduct Double-level Data Filtering (DDF), which includes source data filtering and labeled data filtering, using both non-EMLLM and EMLLM for filtering. Cas-SEAT and DDF work together to improve the performance of EMLLMs. Experiments show that Cas-SEAT achieves an average improvement of 22.16% across multiple datasets, and DDF significantly reduces the resource consumption of training  [^1] The code and a portion of the Cas-SEAT dataset are available at https://github.com/HelloZicky/Cas-SEAT.
Multimodal LLM, Self-evaluation, Chain-of-thought
1. Introduction
Figure 1. (a) A sample from the dataset used for training and inference of multimodal large language models, containing images, questions, and answers. (b) Overview of Chain-of-Thought (CoT) reasoning, self-evaluation reasoning, and their corresponding enhancement methods. (c) The proposed computational method, Cas-SEAT. (d) The proposed dataset construction method, DDF, which provides the Cas-SEAT Dataset for Cas-SEAT. (e) Comparison of CoT reasoning ability, self-evaluation ability, and overall performance. Symbols “–”, “↑”, “↓”, “↑↑”, and “↓↓” indicate comparable, improved, degraded, significantly improved, and significantly degraded performance, respectively.
@@ 1. Introduction
In recent years, multimodal large language models (MLLMs) (Liu et al., 2024a; Wang et al., 2024a) have developed rapidly. However, the growing demand for training and deployment in resource-constrained environments such as universities, hospitals, and communities has spurred interest in developing efficient multimodal large language models (EMLLMs) with smaller parameter sizes, such as 7B, 2B, or even 1B. Thanks to advances in model architecture, training methods, and techniques like chain-of-thought (CoT) reasoning (Wei et al., 2022c), EMLLMs can now generate step-by-step reasoning processes that are more logical and coherent.
With the widespread use of CoT reasoning, the accuracy of each step in the reasoning process has received increasing attention (Madaan et al., 2024; Zheng et al., 2023; Liu et al., 2023b), which we refer to as self-evaluation. However, directly applying these techniques to EMLLMs has not been effective (as shown by the black bars in Figure 3, where the improvement is minimal). This is primarily due to two reasons: (1) EMLLMs tend to “cut corners” when answering questions, making the reasoning process very short. (2) Due to the limited parameter size of EMLLMs, their self-evaluation ability is weak unless enhanced training is performed, rather than just pretraining. More critically, before deploying EMLLMs in applications, traditional supervised fine-tuning (SFT) is often applied to downstream tasks, which further weakens their already fragile self-evaluation ability (as shown in Table 2, where fine-tuned EMLLMs perform worse in self-evaluation (Line 14) compared to pretrained models (Line 13)).
To address these challenges, we intuitively propose a Self-Evaluation Augmented Training (SEAT) method to enhance the self-evaluation capability of EMLLMs. SEAT leverages more powerful EMLLMs to perform CoT reasoning and evaluate them. By optimizing these evaluations, we generate high-quality reasoning samples that help improve the overall reasoning quality and self-evaluation ability of the target model (referred to as the “main model” in this paper). However, the vanilla SEAT method also faces two challenges: (1) Merging reasoning and evaluation tasks into a single prompt leads to excessively long inputs and outputs. From the input perspective, a long prompt is required for EMLLMs to handle both step-by-step reasoning and self-evaluation. From the output perspective, EMLLMs need to perform step-by-step reasoning followed by step-by-step self-evaluation. However, EMLLMs struggle with handling long token inputs and outputs. (2) The CoT reasoning ability of EMLLMs significantly drops after adding self-evaluation data for enhanced training, and CoT reasoning is the foundation of self-evaluation (as shown in the SEAT reasoning and evaluation performance in Tables 1, 2, 3). To overcome these limitations, we further improved SEAT by proposing the Cascaded Self-Evaluation Augmented Training (Cas-SEAT) method. Cas-SEAT decouples the reasoning and self-evaluation processes into independent tasks, using a cascade of brief prompts, with each prompt focusing on a specific task. In order for Cas-SEAT to work, we primarily use labeled CoT data, mixed with a small amount of self-evaluation data. The mixed data is used for training the main model to ensure that the CoT reasoning ability is preserved while enhancing the self-evaluation capability of EMLLMs. Due to the weaker generalization ability of EMLLMs, many data beneficial for training large MLLMs are not suitable for EMLLMs, so we use Double-level Data Filtering (DDF) for data selection. DDF includes source data filtering and labeled data filtering, and the filtering methods involve both non-EMLLM and EMLLM filtering. DDF primarily considers selecting data beneficial for EMLLMs training from six aspects: image quality, text quality, text length, text format, problem domain, and problem difficulty. This ensures the model’s responses are standardized while avoiding data that exceeds the capabilities of EMLLMs. DDF significantly reduces the amount of training data for Cas-SEAT, lowering its training cost and ensuring its applicability in resource-constrained environments. As the first work to explore self-evaluation augmented training, Cas-SEAT and the Cas-SEAT-DDF dataset we constructed provide a valuable foundation and reference for future research in this field.
Experimental results show that Cas-SEAT improves self-evaluation ability by 19.68%, 55.57%, and 46.79% on the MathVista, Math-V, and We-Math datasets, respectively, significantly outperforming existing methods. Furthermore, with the combined effect of Cas-SEAT and DDF, we can use a 7B parameter open-source EMLLMs for data labeling, and the 7B model’s performance exceeds that of a 13B model, fully demonstrating the potential of our method.
@@
  • We analyzed the bottlenecks limiting the self-evaluation ability of EMLLMs and proposed the SEAT method to enhance this ability. To our knowledge, this is the first research focused on self-evaluation for EMLLMs.
@@
  • We designed Cas-SEAT, which effectively addresses the challenges EMLLMs face when handling lengthy prompt inputs and long CoT and self-evaluation outputs. It successfully improves self-evaluation ability while maintaining CoT reasoning capability.
@@
  • We designed DDF and constructed the Cas-SEAT-DDF dataset, the first dataset tailored for self-evaluation augmented training of EMLLMs. It is cost-effective and efficient, facilitating future research.
  • We conducted extensive experiments on multiple datasets, exploring the applicability of our method across different model architectures and parameter sizes. The results show that our method not only significantly improves the performance of EMLLMs but also outperforms many models that are much larger than the main model.
2. Related Work
@@
CoT Reasoning. Chain-of-Thought (CoT) reasoning (Huang et al., 2022; Hoffmann et al., 2022; Chowdhery et al., 2022; Wei et al., 2022a; Zhang et al., 2022; Kojima et al., 2022; Wei et al., 2022b) improves the performance of large language models (LLMs) by forcing them to perform step-by-step reasoning. Subsequent improvements in CoT reasoning for LLMs have been made in areas such as zero-shot CoT prompting (Kojima et al., 2022), few-shot CoT prompting (Wei et al., 2022b; Zhang et al., 2022), self-consistency (Wang et al., 2022b), multiple reasoning paths Wang et al. (2022b, a), minimal-to-maximal prompting (Zhou et al., 2022), dynamic minimal-to-maximal prompting (Drozdov et al., 2022), guided training (Zelikman et al., 2022), and self-training (Zelikman et al., 2022; Huang et al., 2022). To optimize CoT prompts, Li et al. (2022a) proposed filtering CoT prompts using a voting classifier before the final prediction, (Shum et al., 2023) focused on prompt enhancement and selection, and meta-heuristic methods (Pan et al., 2023) and meta-graph prompts (Pan et al., 2024) have also been employed to further optimize CoT prompting. Existing CoT research paid little attention to correctness checking at each step, particularly lacking focus on EMLLMs. However, for EMLLMs, CoT reasoning is more prone to shortcuts and mistakes. Our Cas-SEAT improves the length and accuracy of CoT reasoning in EMLLMs through self-evaluation.
LLM Self-Evaluation. Using additional evaluators to assess the correctness of inference data has been proven effective (Li et al., 2022b; Xu et al., 2024; Liu et al., 2023a), but self-evaluation eliminates the need for additional annotations and evaluators (Li et al., 2022a), making it more efficient. Currently, LLMs demonstrate strong calibration capabilities, and more and more research focuses on using prompts to enable LLMs to perform self-evaluation (Rae et al., 2021; Kadavath et al., 2022; Guo et al., 2017; Kadavath et al., 2022; Jiang et al., 2021; Kuhn et al., 2023; Fu et al., 2023; Golovneva et al., 2022; Zhang et al., 2023; Shinn et al., 2023; Madaan et al., 2024; Paul et al., 2023). Kocmi and Federmann (Kocmi and Federmann, 2023) and Xu et al. (Xu et al., 2023) designed methods to guide LLMs to generate more fine-grained corrective annotations. In addition, Koo et al. (Koo et al., 2023), Zheng et al. (Zheng et al., 2023), Chang et al. (Chang et al., 2023), Deutsch et al. (Deutsch et al., 2022), and Liu et al. (Liu et al., 2023b) explored issues of self-amplifying bias and fairness in LLMs. Notably, the scaling up of models plays a crucial role in improving calibration capabilities (Rae et al., 2021; Wei et al., 2022a). Compared to larger MLLMs, EMLLMs are fragile, with inherently weaker self-evaluation abilities and an inability to handle long token inputs and outputs. Moreover, the foundation of self-evaluation—the CoT reasoning ability—is also weak. Our Cas-SEAT focuses on these issues and significantly improves the self-evaluation ability of EMLLMs.
3. Methodology
@@
Figure 2. Overview of the method. It illustrates the prompts designed for Augmented Self-Evaluation and Augmented Cascading Self-Evaluation, along with the corresponding training data generated.
3.1. Problem Formulation and Notations
3.1.1. Data and Model.
@@ 3.1. Problem Formulation and Notations
The primary EMLLM model is denoted as $\mathcal{M}$, with parameters $\Theta_{0}$. To achieve better performance on the target task, $\mathcal{M}$ needs to be fine-tuned on the dataset $\mathcal{D}$, We use $\{\mathcal{X},\mathcal{Y}\}$ to represent a sample, $\mathcal{X}$ include images $\mathcal{I}$ and queries $\mathcal{Q}$, and $\mathcal{Y}$ represents answers. After fine-tuning, $\Theta_{0}$ will be optimized to $\Theta$. The output of $\mathcal{M}$ are denoted as $\hat{\mathcal{Y}}$. $\mathcal{D}_{\rm{se}}$ represents the self-evaluation data generated by $\mathcal{M}$. A more powerful EMLLM model is denoted as $\mathcal{G}$, which generates CoT reasoning data based on $\mathcal{D}$, denoted as $\mathcal{S}_{\rm{cot}}$. After evaluating $\mathcal{S}_{\rm{cot}}$, $\mathcal{G}$ produces self-evaluation training data, denoted as $\mathcal{S}_{\rm{se}}$. A small subset of $\mathcal{S}_{\rm{cot}}$ is selected, denoted as $\mathcal{S}_{\rm{cot}}^{\rm{sub}}$, and after being evaluated by $\mathcal{G}$, generates cascaded self-evaluation training data, denoted as $\mathcal{S}_{\rm{cse}}$.
3.1.2. Prompt.
@@
The prompts used for $\mathcal{M}$ inference and self-evaluation are denoted as $P$ and $P_{se}$, respectively. The prompts used by $\mathcal{G}$ to generate cascading CoT data and evaluation data are denoted as $P_{\rm{cot}}$ and $P_{se}$, respectively. The prompts used by $\mathcal{G}$ to generate CoT data and evaluation data are denoted as $P_{\rm{cot}}$ and $P_{\rm{{cse}}}$, respectively. Identical symbols indicate identical prompts.
3.1.3. Formula.
We use $l$ to represent the loss function and $\mathcal{L}$ to represent the total loss.
@@
3.2. Double-level Data Filtering
Our data filtering process is conducted in two stages: source data filtering and labeled data filtering. The filtering methods include EMLLM filtering and non-EMLLM filtering. Samples that are filtered out are collectively referred to as “rejected samples” in this paper. In the source data filtering stage, we use the MathV360K dataset (Shi et al., 2024), a commonly used dataset for multimodal large language model training, and apply EMLLM filtering. The filtering criteria are as follows: 1) Image quality: Based on image clarity. Samples with low image clarity are discarded. 2) Text quality: Based on the relevance between text and image and whether the question is clear and well-defined. Samples with mismatched text and images, or with vague questions, are discarded. 3) Question domain: Based on whether specialized domain knowledge is required. Overly specialized scientific questions, such as medical CT image diagnostics, are removed. 4) Question difficulty: Based on whether a more powerful model can answer it correctly. We randomly sample some samples and use Qwen2-VL-7B to attempt answers. If the accuracy of the answers is close to random values, the sample is discarded. In the labeled data filtering stage, we use Qwen2-VL-7B to label the data, followed by non-EMLLM filtering based on Qwen’s responses. The filtering criteria are as follows: 1) Text quality: Samples with garbled responses or responses in languages other than English are discarded. 2) Text length: Samples with excessively long responses are discarded. 3) Text format: The response format is standardized, and samples that do not conform to the format are discarded.
3.3. Vanilla SEAT
Figure 3. Comparison of evaluation in terms of model performance improvement